OpenStudy (anonymous):

Solve the problem. During the 1998-1999 Junior Hockey League season, the Sharks played 54 games. Together, their wins and loses totaled 46. They tied 5 fewer games than they lost. How many games did they win that season? A. 8 games B. 33 games C. 13 games D. 29 games

OpenStudy (anonymous):

Lets set up some variables w +l = 46 t = l-5 w+l+t = 54

OpenStudy (anonymous):

by interchanging the first equation into the last one - you should be able to determine value of T

OpenStudy (anonymous):

from that - determine value of L with 2nd equation and last, determine W value with 1st equation. Please tell me what you got for T L and W

OpenStudy (anonymous):

is it 46?

OpenStudy (whpalmer4):

Let W be the number of games won, L be the number lost, and T be the number tied. They played 54 games, so W + L + T = 54. The count of wins and losses was 46, so W+L = 46. This means we can write W+L+T = 54 (W+L) + T = 54 46 + T = 54 T = 8 They tied 5 few games than they lost: T+5 = L 8+5 = L = 13 W+L = 46 W+13 = 46 W = ?